You own a portfolio that has $3,140 invested in Stock A and $4,300 invested in Stock B. Assume the expected returns on these stocks are 9 percent and 14 percent, respectively. What is the expected return on the portfolio? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1

Total value = 3,140 + 4,300 = 7,440

Expected return = (3,140 / 7,440)*0.09 + (4,300 / 7,440)*0.14

Expected return = 0.037984 + 0.080914

Expected return = 0.1189 or 11.89%